# Diagnosing crosstalk in large-scale QPUs using zero-entropy classical shadows

### v2

@@ II.4 Routing application
The entanglement entropy, $S_{ij}$, of non-local pairs of qubits can be used to detect leakage of information in a non-local way. It is believed that this non-local crosstalk in at least some cases can occur at the multiplexing readout stage [29, 30]. To determine if this information gained through ZECS provides a better strategy for routing qubits in the IBM Heavy-Hex topology than the method used by the qiskit transpiler, the LR-QAOA is used. LR-QAOA is employed in combinatorial optimization applications and is an approximation of a digital adiabatic protocol consisting of a fixed set of linear annealing parameters in the QAOA [31, 32, 28, 33, 34].
@@
For routing, qiskit uses an algorithm called VF2Postlayout [35] that consists of finding the best isomorphic subgraphs to the input circuit using a heuristic objective function derived from the calibration data of the device. In qiskit, there are four levels of routing optimization (from 0 to 3), with 3 being the level where the most effort in the algorithm occurs to find the layout with the lowest error.
@@
To compare both methods, a weighted maxcut problem (WMC) is used with weights randomly selected from the options [0.1,0.2,0.5,2.0] on a 20-qubit 1D-chain topology. LR-QAOA is run from $p=3$ to 100 layers with a $\Delta_{\gamma,\beta}=1.0$ (see Eq. (4) in [28]). The ZECS and qiskit methodologies are run consecutively on ibm_brisbane with 1,000 shots. The performance is calculated using the approximation ratio (see Eq. (11) in [28]).
@@ III Results
Fig. 3 shows a comparison of the infidelity versus the number of snapshots resulting from the CS and ZECS reconstruction of the density state operator for a 3-qubit EfficientSU2 protocol with 7 repetitions. Two different samplers are used: a noiseless sampler (qasm_simulator) and ibm_lagos to show the effect of noise on the reconstruction. Under the assumption of a noiseless procedure (green lines), at $N=1,000$ snapshots, the qasm_simulator reaches an infidelity of $\approx 0.001$ for both CS and ZECS. In the case of a real device (orange curves), using CS, the infidelity stabilizes at around $0.25$ while using ZECS, more of the noise is removed, obtaining an infidelity of $0.057$.
In terms of fidelity, at the end of 10,000 snapshots, CS reaches a $F=0.756$ while ZECS reaches a $F=0.943$. ZECS is an assumption-free methodology and shows the information that a QPU can replicate from a quantum protocol. This is an advantage for diagnosis and characterization, allowing the focus to be on the inherent noise of the device instead of that associated with the instability of the device and the number of snapshots. An extended study of the noise on ibm_lagos is presented in Appendix A.3.
@@
Figure  4(a) shows the relation between the infidelity, $1-F$, and the trace distance, $D$ (See Sec.A.1), for two experiments, 1 and 10 layers of EfficientSU2 in pairs of qubits of ibm_brisbane using CS and ZECS. In Table 2, there is a list of the qubits involved in the EfficientSU2 experiments. In both metrics, a value of zero is desired, and the ZECS can correct most of the noise, bringing the error to an isentropic line where $1-F$ and $|D|$ are correlated. In contrast, in the CS case, because of an unstable reconstruction of the density state operator, there is no strong correlation between the two metrics. Figure  4(b) shows the singular values magnitude for the bipartite systems involved in the $n=10$ EfficientSU2 experiments. The order of them is the same as shown in Table 2. Note that there is a good separation of $\lambda_{1}$ to the other singular values. This is a good indication that noise is not strong enough to forbid a good reconstruction of the density state operator.
Figure 4: (a) ($1-F$) vs. $D$ using CS and ZECS for the 2-qubit experiment of the EfficientSU2 gates for $n=1$ and 10 repetitions on ibm_brisbane. The inset circuit is the EfficientSU2 gate repeated n times with random parameters $\theta_{i}$. Every marker represents a pair of qubit results. (b) Singular values of $\rho_{cs}$ for the bipartite subsystems in the $n=10$ EfficientSU2 experiment.
@@
Figure  5 shows the fidelity information of ibm_brisbane obtained from the reconstruction of density state operators using $N=6,000$ snapshots and the ZECS methodology for 10 repetitions of the EfficientSU2 circuit. Fig 5(a) shows the layout of ibm_brisbane, with edge color representing the fidelity of the density state operator reconstructed. This reconstruction is based on the information of 2, 3, and 4-qubit density state operators. In Table 2, there is a list of the qubits involved in each reconstruction. This fidelity map helps to identify regions where poor performance is present. For example, qubit 68 and its neighboring nodes have the lowest performance. If we compare this fidelity against, for example, the RB fidelity, we can expect a more detailed noise characterization. Because the characterization of 2-qubit gates using RB uses only one expectation value.
Fig.5(b) shows the infidelity of the 2-qubit systems involved in the EfficientSU2 protocol. There is a significant impact of applying the ZECS methodology to the CS information, reducing the infidelity from 0.19(±$\pm 0.08$) to $\pm 0.03$) on average. Furthermore, there is a correlation between the infidelity of CS and ZECS, but not in all cases. For instance, subsystem (27,28) has a lower infidelity than subsystem (26,25) using the CS reconstruction, but once ZECS is applied, subsystem (26,25) has a lower infidelity. This suggests that some kinds of errors are more prone to be corrected. ZECS shows improvements in some cases of 1 order of magnitude in the infidelity compared to CS. It is interesting to note that ZECS is based completely on the CS information and no further assumptions are made. Therefore, the extra information recovered in ZECS comes completely from the QPU. Figs. 5(c) and (d) show the infidelity for the 3-qubit and 4-qubit cases. The recovery of ZECS is also important in these cases. A detriment in fidelity from the 3 and 4-qubit cases when compared with the 2-qubit case results primarily from the dimensionality of the density matrix that is reconstructed. In the case of the 2-qubit problem, there are only $42^{N_{q}}=$ 4 complex parameters, while for 3 qubits, there are 8 and 4 qubits 16.
@@
Now, based on the information provided by Figs. 5 and 6, the best 20 qubits that form a 1D chain are chosen and compared to the best routing procedure of Qiskit. In the inset of Fig. 7(a), the chain chosen using ZECS information is indicated in orange, while the chain chosen by the qiskit transpiler with an optimization level 3 is shown in blue. The routing procedure of qiskit uses calibration information of the device, including the native 2-qubit gates error. In the case of ZECS, the fidelity and entanglement entropy information are used to guide the routing.
@@
$\mathrm{ECR}^{avg}_{error}$ $0.01\ (\pm 4.8\times 10^{-3})$ $0.006\ (\pm 1.8\times 10^{-3})$
$\bar{F}$ $0.983\ (\pm 0.011)$ $0.974\ (\pm 0.015)$
$\bar{S_{ij}}$ $0.299\ (\pm 0.09)$ $0.375\ (\pm 0.155)$
Figure 7(a) shows the results of the optimization solutions using the LR-QAOA of Sec. II.4 for the WMC. The y-axis represents the approximation ratio vs. the number of QAOA layers. This metric indicates the proximity of the QPU’s outputs to the optimal solution of the problem, with $r=1$ indicating a 100% probability of finding the optimal solution of WMC using LR-QAOA. In an ideal case, as $p$ grows, $r$ approaches 1. However, when noise is involved, $r$ grows to the point where the LR-QAOA is stronger than the noise inherent in the device. As can be seen, using the ZECS routing, the chosen set of qubits improves the lifetime of the LR-QAOA compared to the routing chosen by qiskit by more than 33%, i.e., from $p=50$ to $p=75$, and the peak of the approximation ratio from $r=0.750$ to $r=0.785$, which is a 10% reduction in error when compared to the ideal approximation ratio of $r=0.876$ at $p=5$. This suggests that the qiskit routing and the two-qubit errors do not fully capture the characteristics needed for a set of qubits to perform well on some tasks that require cohesion between qubits.
@@ IV Conclusions
Experimental validation of the ZECS methodology is performed using ibm_lagos, ibm_brisbane, and ibm_fez. 6,000 shots are demonstrated to be sufficient to recover the information of 2, 3, and 4 qubit density state operators. Using this information, a diagnostic in terms of fidelity, trace distance, and entanglement entropy is given. These metrics show how pairs of qubits perform when they are involved in gate operations. Entanglement entropy provides additional information about the possible leakage of information or crosstalk between pairs of qubits, idle qubits, and two neighboring pairs of qubits.
Using ZECS, regions where quantum devices have low performance are identified and can, thus, be avoided. This methodology is shown to have better performance in finding a set of qubits for a quantum optimization application compared to the qiskit routing procedure, improving the lifetime of the quantum algorithm by more than 33% and the approximation ratio from $r=0.750$ to $r=0.785$. The improvement is the result of a more informed decision. While qiskit routes are based on the collected calibration information of disjoint terms, routing using ZECS uses detailed information from the density state operator, like fidelity and entanglement entropy. These are meaningful metrics that can be correlated to noise and crosstalk. This reconstruction is flexible, i.e., the qubits for the reconstruction can be chosen arbitrarily, although it is necessary to be aware that the number of snapshots needed grows exponentially with the number of qubits involved in the reconstruction. However, with only 6000 shots, a good estimation of the density state operator for up to 4 qubits can be given.
The density state operator reconstruction has the advantage that one can unveil hidden noise elements like unwanted correlations using the entanglement entropy. This characteristic is tested to uncover non-local crosstalk that can explain partially why some pairs of qubits have low performance. For example, reconstructing the density state operator of nonconnected pairs of qubits (2,3) and (19,20) on ibm_brisbane shows an entanglement entropy that deviates considerably from the mean entanglement entropy calculated for the interaction of (19,20) with all other pairs. It suggests that there is a leakage of information that can occur at the multiplexing readout stage.
@@ A.1 Metrics
where $\rho_{1}$ is the density state reconstruction and $\rho_{2}$ the ideal density state operator. The trace distance [36] is defined by
@@
The concurrence [37] is a metric of the entanglement of pairs of qubits and is given by
@@
The entanglement entropy is expressed as
$S_{ab}=-\rho_{a}\log(\rho_{a})=-\rho_{b}\log(\rho_{b}),$ (14)
where $\rho_{a}={\rm Tr}_{b}(\rho)$. When the systems are disentangled, i.e., $\rho=\rho_{a}\otimes\rho_{b}$, $S_{ab}=0$.
@@ A.2 Zero Entropy
The ZE methodology is general and can be applied to mixed states. In Fig. 9, the capabilities of the ZE methodology to recover information from a noisy density state operator are shown. Random perturbations on a Bell state, $|\psi\rangle=\frac{1}{\sqrt{2}}\left(|00\rangle+|11\rangle\right)$ are used and information is recovered using ZE. The random perturbation presented in [23] is given by
@@
The density operator $\rho_{ze}$ is first constructed from $\rho$ using Eqs. 7 and 8. In Fig. 9, $\rho$ and $\rho_{ze}$ are compared in terms of the (a) infidelity ($1-F$), (b) $D$, and (c) $C$. These metrics are intended to give a sense of how close both states are to $\rho_{0}$. For $1-F$ and $D$, it is seen that for perturbations with $\sigma<0.3$, the $\rho_{ze}$ is closer to the ideal $\rho_{0}$. In practice, this means that for QPUs with low error rates, ZE allows the reconstruction of a closer representation of $\rho_{0}$. In terms of $C$, which measures entanglement in pairs of qubits, using ZE, the entanglement remains stronger and for longer than the noisy case.
@@
Figure 10 shows the 4 eigenvalues of the Bell state for the $ibm\_brisbane$ experiment (Fig. 4(b)) when random perturbations are applied. As can be seen, as the strength of $\sigma$ increases, the information of the first eigenvalue decreases and the other eigenvalues begin to become more relevant. That the curves never cross as the strength of the noise increases indicates that the first eigenvalue can always be the largest of all the eigenvalues.
@@
Fig.11(b) and (c) show the real (blue) and imaginary (magenta) components of $\rho$ at $N=10,000$. Because of noise and limitations in sampling, the results deviate from the noiseless result. However, after applying ZECS to the CS information, part of the noise is removed so that the real and imaginary parts of the reconstructed density operator approach those for the noiseless simulation.

### v3

@@ II.5 Routing application
The entanglement entropy, $S_{ij}$, of non-local pairs of qubits can be used to detect leakage of information in a non-local way. It is believed that this non-local crosstalk in at least some cases can occur at the multiplexing readout stage [Jurcevic2021, Jerger2012]. To determine if this information gained through ZECS provides a better strategy for routing qubits in the IBM Heavy-Hex topology than the method used by the qiskit transpiler, the LR-QAOA is used. LR-QAOA is employed in combinatorial optimization applications and is an approximation of a digital adiabatic protocol consisting of a fixed set of linear annealing parameters in the QAOA [Farhi2014, Willsch2022, montanezbarrera2024universal, hess2024effectiveembeddingintegerlinear, Kremenetski2023].
@@
For routing, qiskit uses an algorithm called VF2Postlayout [Nation_2023] that consists of finding the best isomorphic subgraphs to the input circuit using a heuristic objective function derived from the calibration data of the device. In qiskit, there are four levels of routing optimization (from 0 to 3), with 3 being the level where the most effort in the algorithm occurs to find the layout with the lowest error.
@@
To compare both methods, a weighted maxcut problem (WMC) is used with weights randomly selected from the options [0.1,0.2,0.5,2.0] on a 20-qubit 1D-chain topology. LR-QAOA is run from $p=3$ to 100 layers with a $\Delta_{\gamma,\beta}=1.0$ (see Eq. (4) in [montanezbarrera2024universal]). The ZECS and qiskit methodologies are run consecutively on ibm_brisbane with 1,000 shots. The performance is calculated using the approximation ratio (see Eq. (11) in [montanezbarrera2024universal]).
@@ III Results
Fig. 4 shows a comparison of the infidelity versus the number of snapshots resulting from the CS and ZECS reconstruction of the density state operator for a 3-qubit EfficientSU2 protocol with 7 repetitions. Two different samplers are used: a noiseless sampler (qasm_simulator) and ibm_lagos to show the effect of noise on the reconstruction. Under the assumption of a noiseless procedure (green lines), at $N=1,000$ snapshots, the qasm_simulator reaches an infidelity of $\approx 0.001$ for both CS and ZECS. In the case of a real device (orange curves), using CS, the infidelity stabilizes at around $0.25$ while using ZECS, more of the noise is removed, obtaining an infidelity of $0.057$. An extended study of the noise on ibm_lagos is presented in Appendix A.3.
@@
Figure  5(a) shows the relation between the infidelity, 1F1-F, and the trace distance, $D$ (See Sec.A.1), for two experiments, 1 and 10 layers of EfficientSU2 in pairs of qubits of ibm_brisbane using CS and ZECS. In Table 2, there is a list of the qubits involved in the EfficientSU2 experiments. In both metrics, a value of zero is desired, and the ZECS can correct most of the noise, bringing the error to an isentropic line where 1F1-F and $|D|$ are correlated. In contrast, in the CS case, because of an unstable reconstruction of the density state operator, there is no strong correlation between the two metrics. Figure  5(b) shows the singular values magnitude for the bipartite systems involved in the $n=10$ EfficientSU2 experiments. The order of them is the same as shown in Table 2. Note that there is a good separation of $\lambda_{1}$ to the other singular values. This is a good indication that noise is not strong enough to forbid a good reconstruction of the density state operator.
Figure 5: (a) (1F1-F) vs. $D$ using CS and ZECS for the 2-qubit experiment of the EfficientSU2 gates for $n=1$ and 10 repetitions on ibm_brisbane. The inset circuit is the EfficientSU2 gate repeated n times with random parameters $\theta_{i}$. Every marker represents a pair of qubit results. (b) Singular values of $\rho_{cs}$ for the bipartite subsystems in the $n=10$ EfficientSU2 experiment.
@@
Figure  6 shows the fidelity information of ibm_brisbane obtained from the reconstruction of density state operators using $N=6,000$ snapshots and the ZECS methodology for 10 repetitions of the EfficientSU2 circuit. Fig 6(a) shows the layout of ibm_brisbane, with edge color representing the fidelity of the density state operator reconstructed. This reconstruction is based on the information of 2, 3, and 4-qubit density state operators. In Table 2, there is a list of the qubits involved in each reconstruction. This fidelity map helps to identify regions where poor performance is present. For example, qubit 68 and its neighboring nodes have the lowest performance. If we compare this fidelity against, for example, the RB fidelity, we can expect a more detailed noise characterization. Because the characterization of 2-qubit gates using RB uses only one expectation value.
Fig.6(b) shows the infidelity of the 2-qubit systems involved in the EfficientSU2 protocol. There is a significant impact of applying the ZECS methodology to the CS information, reducing the infidelity from 0.19($\pm 0.08$) to 0.03($\pm 0.03$) on average. Furthermore, there is a correlation between the infidelity of CS and ZECS, but not in all cases. For instance, subsystem (27,28) has a lower infidelity than subsystem (26,25) using the CS reconstruction, but once ZECS is applied, subsystem (26,25) has a lower infidelity. This suggests that some kinds of errors are more prone to be corrected. ZECS shows improvements in some cases of 1 order of magnitude in the infidelity compared to CS. It is interesting to note that ZECS is based completely on the CS information and no further assumptions are made. Therefore, the extra information recovered in ZECS comes completely from the QPU. Figs. 6(c) and (d) show the infidelity for the 3-qubit and 4-qubit cases. The recovery of ZECS is also important in these cases. A detriment in fidelity from the 3 and 4-qubit cases when compared with the 2-qubit case results primarily from the dimensionality of the density matrix that is reconstructed. In the case of the 2-qubit problem, there are only $2^{N_{q}}=4$ complex parameters, while for 3 qubits, there are 8 and 4 qubits 16.
@@
Now, based on the information provided by Figs. 6 and 7, the best 20 qubits that form a 1D chain are chosen and compared to the best routing procedure of Qiskit. In the inset of Fig. 8(a), the chain chosen using ZECS information is indicated in orange, while the chain chosen by the qiskit transpiler with an optimization level 3 is shown in blue. The routing procedure of qiskit uses calibration information of the device, including the native 2-qubit gates error. In the case of ZECS, the fidelity and entanglement entropy information are used to guide the routing.
@@
[…]
$\bar{F}$ 0.983(±0.011)0.983\ ($\pm$ 0.011) 0.974(±0.015)0.974\ ($\pm$ 0.015)
$\bar{S_{ij}}$ 0.299(±0.09)0.299\ ($\pm$ 0.09) 0.375(±0.155)0.375\ ($\pm$ 0.155)
Figure 8(a) shows the results of the optimization solutions using the LR-QAOA of Sec. II.5 for the WMC. The y-axis represents the approximation ratio vs. the number of QAOA layers. This metric indicates the proximity of the QPU’s outputs to the optimal solution of the problem, with $r=1$ indicating a 100% probability of finding the optimal solution of WMC using LR-QAOA. In an ideal case, as $p$ grows, $r$ approaches 1. However, when noise is involved, $r$ grows to the point where the LR-QAOA is stronger than the noise inherent in the device. As can be seen, using the ZECS routing, the chosen set of qubits improves the lifetime of the LR-QAOA compared to the routing chosen by qiskit by more than 33%, i.e., from $p=50$ to $p=75$, and the peak of the approximation ratio from $r=0.750$ to $r=0.785$, which is a 10% reduction in error when compared to the ideal approximation ratio of $r=0.876$ at $p=5$. This suggests that the qiskit routing and the two-qubit errors do not fully capture the characteristics needed for a set of qubits to perform well on some tasks that require cohesion between qubits.
@@
III.4 Cross-platform comparison
Figure 10 shows the comparison between ionq_forte, ibm_fez, and a noiseless simulator. From the previous experiment on ibm_fez, result reconstructions with only 1000 samples are used in 60 pairs out of 176 possible pairs, and the same protocol is run on ionq_forte using its 36 qubits (18 pairs) and on the noiseless simulator using 20 qubits (10 pairs). The reconstruction is then made in pairs of 2-qubit subsystems throughout all possible pairs, which grows as $N_{total}=N_{pairs}(N_{pairs}-1)/2$. In the case of ibm_fez, there are 1770 pairs, on ionq_forte 153 pairs, and on the noiseless simulator 45 pairs.
Figure 10(a) presents the entanglement entropy versus fidelity for the three scenarios. Among them, ibm_fez shows the strongest impact of crosstalk, affecting many qubit pairs. Note that some part of the entanglement entropy can still be attributed to sampling limitations since even in the noiseless simulation, a residual $S_{ab}$ remains.
Figure 10(b) summarize the $S_{ab}$ of the different platforms. The median $S_{ab}$ values (with the upper whiskers in parentheses) are $[0.22(0.46),0.16(0.28),0.12(0.18)]$ for the ibm_fez, ionq_forte, and noiseless case, respectively. Furthermore, ibm_fez shows 190 outliers, compared with only 2 outliers in the other two cases. These findings are consistent with previously reported performance across both platforms when running the same quantum applications [montanezbarrera2024universal, montanezbarrera2025evaluating]. Despite ibm_fez having a lower average two-qubit gate error rate ($2.6\times 10^{-3}$) than ionq_forte ($10.2\times 10^{-3}$), the performance of the ionq_forte in these applications is better.
Figure 10: Cross-platform comparison of ZECS between trapped ions and superconducting QPUs of 2-qubit circuit using 1000 samples. (a) Fidelity versus entanglement entropy. Markers represent all the 2-subsystem pairs of each device. (b) The box shows the interquartile range (IQR) for the entanglement entropy, with the central red line indicating the median. Whiskers extend to data within 1.5IQR, caps mark their endpoints, and circles are outliers.
@@ IV Conclusions
Experimental validation of the ZECS methodology is performed using ibm_lagos, ibm_brisbane, ibm_fez, and ionq_forte. 1,000 shots are demonstrated to be sufficient to recover the information of 2, 3, and 4 qubit density state operators. Using this information, a diagnostic in terms of fidelity, trace distance, and entanglement entropy is given. These metrics show how pairs of qubits perform when they are involved in gate operations. Entanglement entropy provides additional information about the crosstalk between pairs of qubits. This is, to our knowledge, the first cross-platform quantification of crosstalk under identical circuit conditions.
Using ZECS, regions where quantum devices have low performance are identified and can, thus, be avoided. This methodology is shown to have better performance in finding a set of qubits for a quantum optimization application compared to the qiskit routing procedure, improving the lifetime of the quantum algorithm by more than 33% and the approximation ratio from $r=0.750$ to $r=0.785$. The improvement is the result of a more informed decision.
While qiskit routes are based on the collected calibration information of disjoint terms, routing using ZECS uses detailed information from the density state operator, like fidelity and entanglement entropy. These are meaningful metrics that can be correlated to noise and crosstalk. This reconstruction is flexible, i.e., the qubits for the reconstruction can be chosen arbitrarily, although it is necessary to be aware that the number of snapshots needed grows exponentially with the number of qubits involved in the reconstruction. However, with only 6000 shots, a good estimation of the density state operator for up to 4 qubits can be given.
We extended the ZECS experiments to ionq_forte, a trapped-ion QPU from IonQ. On this platform, crosstalk is present but less pronounced than on ibm_fez. The results also reveal that certain subsystems of ibm_fez are highly correlated, consistent with the large number of outliers observed. This protocol can also be extended to other platforms, such as neutral-atom and photonic-based QPUs. The only requirement is the availability of single-qubit gates that include at least $R_{X}$ and $R_{Z}$ rotations.
The density state operator reconstruction has the advantage that one can unveil hidden noise elements like unwanted correlations using the entanglement entropy. This characteristic is tested to uncover non-local crosstalk that can explain partially why some pairs of qubits have low performance. For example, reconstructing the density state operator of nonconnected pairs of qubits (2,3) and (19,20) on ibm_brisbane shows an entanglement entropy that deviates considerably from the mean entanglement entropy calculated for the interaction of (19,20) with all other pairs. It suggests that information leakage may occur at the multiplexed readout stage, a phenomenon that has been previously reported in [Heinsoo_2018, bakr2024multiplexed, Maurya_2024].
@@ A.1 Metrics
where $\rho_{1}$ is the density state reconstruction and $\rho_{2}$ the ideal density state operator. The trace distance [Nielsen2011] is defined by
@@
The concurrence [Hill1997] is a metric of the entanglement of pairs of qubits and is given by
@@ A.2 Zero Entropy
The ZE methodology is general and can be applied to mixed states. In Fig. 11, the capabilities of the ZE methodology to recover information from a noisy density state operator are shown. Random perturbations on a Bell state, $|\psi\rangle=\frac{1}{\sqrt{2}}\left(|00\rangle+|11\rangle\right)$ are used and information is recovered using ZE. The random perturbation presented in [Montanez-Barrera2022] is given by
@@
The density operator $\rho_{ze}$ is first constructed from $\rho$ using Eqs. 7 and 8. In Fig. 11, $\rho$ and $\rho_{ze}$ are compared in terms of the (a) infidelity (1F1-F), (b) $D$, and (c) $C$. These metrics are intended to give a sense of how close both states are to $\rho_{0}$. For 1F1-F and $D$, it is seen that for perturbations with $\sigma<0.3$, the $\rho_{ze}$ is closer to the ideal $\rho_{0}$. In practice, this means that for QPUs with low error rates, ZE allows the reconstruction of a closer representation of $\rho_{0}$. In terms of $C$, which measures entanglement in pairs of qubits, using ZE, the entanglement remains stronger and for longer than the noisy case.
@@
Figure 12 shows the 4 eigenvalues of the Bell state for the ibm_brisbaneibm\_brisbane experiment (Fig. 5(b)) when random perturbations are applied. As can be seen, as the strength of $\sigma$ increases, the information of the first eigenvalue decreases and the other eigenvalues begin to become more relevant. That the curves never cross as the strength of the noise increases indicates that the first eigenvalue can always be the largest of all the eigenvalues.
@@
Fig.13(b) and (c) show the real (blue) and imaginary (magenta) components of $\rho$ at $N=10,000$. Because of noise and limitations in sampling, the results deviate from the noiseless result. However, after applying ZECS to the CS information, part of the noise is removed so that the real and imaginary parts of the reconstructed density operator approach those for the noiseless simulation. In terms of fidelity, at the end of 10,000 snapshots, CS reaches a $F=0.756$ while ZECS reaches a $F=0.943$. ZECS is an assumption-free methodology and shows the information that a QPU can replicate from a quantum protocol. This is an advantage for diagnosis and characterization, allowing the focus to be on the inherent noise of the device instead of that associated with the instability of the device and the number of snapshots.